# Restricted $L_{\infty}$-algebras and a derived Milnor-Moore theorem

### v1

@@ 2.4. Truncating $\infty$-operads
Let ${\mathcal{C}}$ be a symmetric monoidal $\infty$-category compatible with the initial object that admits small colimits. The embedding $\kappa:\Sigma_{\geq{\mathrm{n}}}\subset\Sigma_{\geq 1}$ induces a colocalization
$\kappa_{!}:{\mathrm{sSeq}}({\mathcal{C}})_{\geq{\mathrm{n}}}\rightleftarrows{\mathrm{sSeq}}({\mathcal{C}})_{\geq 1}:\kappa^{\ast},$
where the right adjoint is restriction and the fully faithful left adjoint extends under degree ${\mathrm{n}}$ by the initial object. By Lemma 2.43 the embedding $\kappa_{!}:{\mathrm{sSeq}}({\mathcal{C}})_{\geq{\mathrm{n}}}\subset{\mathrm{sSeq}}({\mathcal{C}})_{\geq 1}$ is ${\mathrm{sSeq}}({\mathcal{C}})_{\geq 1}$-linear. Thus for every non-unital $\infty$-operad ${\mathcal{O}}$ in ${\mathcal{C}}$ we obtain a colocalization
${\mathrm{RMod}}_{{\mathcal{O}}}({\mathrm{sSeq}}({\mathcal{C}})_{\geq{\mathrm{n}}})\rightleftarrows{\mathrm{RMod}}_{{\mathcal{O}}}({\mathrm{sSeq}}({\mathcal{C}})_{\geq 1}):\tau_{\geq{\mathrm{n}}}.$
@@
Let ${\mathcal{C}}$ be a presentably symmetric monoidal $\infty$-category and ${\mathcal{O}}$ a non-unital $\infty$-operad in ${\mathcal{C}}$ such that ${\mathcal{O}}_{1}={\mathbb{1}}$. Let ${\mathrm{n}}\geq 1$ and ${\mathrm{X}}$ a right ${\mathcal{O}}$-module in ${\mathcal{C}}$ concentrated in degree ${\mathrm{n}}$. The space of right ${\mathcal{O}}$-module structures on ${\mathrm{X}}$ is contractible.
@@
As ${\mathrm{sSeq}}({\mathcal{C}})_{\geq 1}$ is presentable, the small colimits preserving functor $(-)\circ{\mathrm{X}}:{\mathrm{sSeq}}({\mathcal{C}})_{\geq 1}\to{\mathrm{sSeq}}({\mathcal{C}})_{\geq 1}$ admits a right adjoint $[{\mathrm{X}},-].$ Thus the space of right ${\mathcal{O}}$-module structures on ${\mathrm{X}}$ is equivalent to the space of maps of $\infty$-operads ${\mathcal{O}}\to[{\mathrm{X}},{\mathrm{X}}]$ in ${\mathcal{C}}.$
For every ${\mathrm{k}}\in\Sigma_{\geq 1}$ the functor ${\mathrm{sSeq}}({\mathcal{C}})_{\geq 1}\to{\mathcal{C}}$ evaluating at ${\mathrm{k}}$ admits a left adjoint ${\mathrm{F}}_{\mathrm{k}}$ sending ${\mathrm{Z}}$ to the symmetric sequence concentrated in degree ${\mathrm{k}}$ with value $\Sigma_{\mathrm{k}}\times{\mathrm{Z}}.$ Hence by adjointness we get the following: If ${\mathrm{k}}>1$, we find that
${\mathcal{C}}({\mathrm{Z}},[{\mathrm{X}},{\mathrm{X}}]_{\mathrm{k}})\simeq{\mathrm{sSeq}}({\mathcal{C}})_{\geq 1}({\mathrm{F}}_{\mathrm{k}}({\mathrm{Z}})\circ{\mathrm{X}},{\mathrm{X}})$
is contractible so that $[{\mathrm{X}},{\mathrm{X}}]_{\mathrm{k}}$ is the final object. Otherwise, we find that
${\mathcal{C}}({\mathrm{Z}},[{\mathrm{X}},{\mathrm{X}}]_{1})\simeq{\mathrm{sSeq}}({\mathcal{C}})_{\geq 1}({\mathrm{F}}_{1}({\mathrm{Z}})\circ{\mathrm{X}},{\mathrm{X}})\simeq{\mathcal{C}}^{\Sigma_{\mathrm{n}}}({\mathrm{Z}}\otimes{\mathrm{X}}_{\mathrm{n}},{\mathrm{X}}_{\mathrm{n}})\simeq{\mathcal{C}}({\mathrm{Z}},[{\mathrm{X}}_{\mathrm{n}},{\mathrm{X}}_{\mathrm{n}}]^{\Sigma_{n}}),$
where $[{\mathrm{X}}_{\mathrm{n}},{\mathrm{X}}_{\mathrm{n}}]$ is the internal hom in ${\mathcal{C}}$, and so $[{\mathrm{X}},{\mathrm{X}}]_{1}\simeq[{\mathrm{X}}_{\mathrm{n}},{\mathrm{X}}_{\mathrm{n}}]^{\Sigma_{n}}$. Using the localization
$(-)_{\leq 1}:{\mathrm{Op}}({\mathcal{C}})^{\mathrm{nu}}\to{\mathrm{Op}}({\mathcal{C}})^{\mathrm{nu}}_{\leq 1}=\mathrm{Alg}({\mathcal{C}})$
the space of maps of $\infty$-operads ${\mathcal{O}}\to[{\mathrm{X}},{\mathrm{X}}]$ in ${\mathcal{C}}$ identifies with the space of maps ${\mathcal{O}}_{1}\to[{\mathrm{X}}_{\mathrm{n}},{\mathrm{X}}_{\mathrm{n}}]^{\Sigma_{n}}$ of associative algebras in ${\mathcal{C}}$. But by assumption ${\mathcal{O}}_{1}\simeq{\mathbb{1}}$ so that ${\mathcal{O}}_{1}$ is the initial associative algebra in ${\mathcal{C}}.$
@@ 2.5. Hopf $\infty$-operads
Let ${\mathcal{C}}$ be a symmetric monoidal $\infty$-category. By [23, Proposition 3.2.4.3.] the $\infty$-category $\mathrm{Alg}_{{\mathbb{E}}_{\infty}}({\mathcal{C}})$ carries a canonical symmetric monoidal structure such that the forgetful functor $\mathrm{Alg}_{{\mathbb{E}}_{\infty}}({\mathcal{C}})\to{\mathcal{C}}$ is symmetric monoidal. So dually, the $\infty$-category
${\mathrm{coAlg}}_{{\mathbb{E}}_{\infty}}({\mathcal{C}}):=\mathrm{Alg}_{{\mathbb{E}}_{\infty}}({\mathcal{C}}^{\mathrm{op}})^{\mathrm{op}}$
carries a symmetric monoidal structure such that the forgetful functor ${\mathrm{coAlg}}_{{\mathbb{E}}_{\infty}}({\mathcal{C}})\to{\mathcal{C}}$ is symmetric monoidal.
If ${\mathcal{C}}$ admits small colimits, ${\mathrm{coAlg}}_{{\mathbb{E}}_{\infty}}({\mathcal{C}})$ admits small colimits and the forgetful functor ${\mathrm{coAlg}}_{{\mathbb{E}}_{\infty}}({\mathcal{C}})\to{\mathcal{C}}$ preserves small colimits [23, Corollary 3.2.2.5.]. So we can make the following definition:
@@
A Hopf $\infty$-operad in ${\mathcal{C}}$ is an $\infty$-operad in ${\mathrm{coAlg}}_{{\mathbb{E}}_{\infty}}({\mathcal{C}})$.
@@
${\mathrm{Op}}_{\mathrm{Hopf}}({\mathcal{C}}):={\mathrm{Op}}({\mathrm{coAlg}}_{{\mathbb{E}}_{\infty}}({\mathcal{C}})),\ {\mathrm{Op}}_{\mathrm{Hopf}}^{\mathrm{nu}}({\mathcal{C}}):={\mathrm{Op}}({\mathrm{coAlg}}_{{\mathbb{E}}_{\infty}}({\mathcal{C}}))^{\mathrm{nu}}.$
The symmetric monoidal forgetful functor ${\mathrm{coAlg}}_{{\mathbb{E}}_{\infty}}({\mathcal{C}})\to{\mathcal{C}}$ yields forgetful functors
${\mathrm{Op}}_{\mathrm{Hopf}}({\mathcal{C}})\to{\mathrm{Op}}({\mathcal{C}}),\ {\mathrm{Op}}^{\mathrm{nu}}_{\mathrm{Hopf}}({\mathcal{C}})\to{\mathrm{Op}}({\mathcal{C}})^{\mathrm{nu}}.$
@@
Since ${\mathrm{coAlg}}_{{\mathbb{E}}_{\infty}}({\mathcal{C}})$ admits a final object lying over the tensor unit of ${\mathcal{C}},$ by Lemma 2.32 the $\infty$-categories ${\mathrm{Op}}_{\mathrm{Hopf}}({\mathcal{C}}),\ {\mathrm{Op}}^{\mathrm{nu}}_{\mathrm{Hopf}}({\mathcal{C}})$ admit a final object lying over the constant symmetric sequence (concentrated in positive degrees) with value the tensor unit.
@@
The non-unital commutative $\infty$-operad $\mathrm{Comm}^{\mathrm{nu}}_{\mathcal{C}}$ in ${\mathcal{C}}$ is the final object of ${\mathrm{Op}}^{{\mathrm{nu}}}_{\mathrm{Hopf}}({\mathcal{C}})$.
@@
The non-counital cocommutative $\infty$-cooperad ${\mathrm{coComm}}^{\mathrm{nu}}_{\mathcal{C}}$ in ${\mathcal{C}}$ is $\mathrm{Comm}^{\mathrm{nu}}_{{\mathcal{C}}^{\mathrm{op}}}$.
@@
${\mathcal{O}}^{\mathrm{nu}}:=\mathrm{Comm}_{\mathcal{C}}^{\mathrm{nu}}\times_{\mathrm{Comm}_{\mathcal{C}}}{\mathcal{O}}$
@@
$\mathrm{Comm}_{\mathcal{C}}^{\mathrm{nu}}\times_{\mathrm{Comm}_{\mathcal{C}}}{\mathcal{O}}$
in ${\mathrm{Op}}({\mathcal{C}})$ exists and is preserved by the forgetful functor ${\mathrm{Op}}({\mathcal{C}})\to{\mathrm{sSeq}}({\mathcal{C}})$ by Lemma 2.32. So it is enough to see that the fiber product $\mathrm{Comm}_{\mathcal{C}}^{\mathrm{nu}}\times_{\mathrm{Comm}_{\mathcal{C}}}{\mathcal{O}}$ in ${\mathrm{Hopf}}({\mathcal{C}})$ exists and is preserved by the forgetful functor ${\mathrm{Hopf}}({\mathcal{C}})\to{\mathrm{sSeq}}({\mathcal{C}})$. Since every fiber product in ${\mathrm{sSeq}}({\mathrm{coAlg}}_{{\mathbb{E}}_{\infty}}({\mathcal{C}}))$ of maps in ${\mathrm{Hopf}}({\mathcal{C}})$ lifts to a fiber product in ${\mathrm{Hopf}}({\mathcal{C}})$ by Lemma 2.32, it remains to prove that the fiber product $\mathrm{Comm}_{\mathcal{C}}^{\mathrm{nu}}\times_{\mathrm{Comm}_{\mathcal{C}}}{\mathcal{O}}$ in ${\mathrm{sSeq}}({\mathrm{coAlg}}_{{\mathbb{E}}_{\infty}}({\mathcal{C}}))$ exists and is preserved by the forgetful functor ${\mathrm{sSeq}}({\mathrm{coAlg}}_{{\mathbb{E}}_{\infty}}({\mathcal{C}}))\to{\mathrm{sSeq}}({\mathcal{C}})$. For that it is enough to show that for every ${\mathrm{n}}\geq 0$ the fiber product $(\mathrm{Comm}_{\mathcal{C}}^{\mathrm{nu}})_{\mathrm{n}}\times_{(\mathrm{Comm}_{\mathcal{C}})_{\mathrm{n}}}{\mathcal{O}}_{\mathrm{n}}$ in ${\mathrm{coAlg}}_{{\mathbb{E}}_{\infty}}({\mathcal{C}})$ exists and is preserved by the forgetful functor ${\mathrm{coAlg}}_{{\mathbb{E}}_{\infty}}({\mathcal{C}})\to{\mathcal{C}}$. This holds by [23, Proposition 3.1.3.3.] if and only if for every $\ell\geq 0$ the morphism
$((\mathrm{Comm}_{\mathcal{C}}^{\mathrm{nu}})_{\mathrm{n}}\times_{(\mathrm{Comm}_{\mathcal{C}})_{\mathrm{n}}}{\mathcal{O}}_{\mathrm{n}})^{\otimes\ell}\to((\mathrm{Comm}_{\mathcal{C}}^{\mathrm{nu}})_{\mathrm{n}})^{\otimes\ell}\times_{((\mathrm{Comm}_{\mathcal{C}})_{\mathrm{n}})^{\otimes\ell}}({\mathcal{O}}_{\mathrm{n}})^{\otimes\ell}$
is an equivalence. The latter map trivially is an equivalence since $\mathrm{Comm}_{\mathcal{C}}^{\mathrm{nu}}\to\mathrm{Comm}_{\mathcal{C}}$ induces an equivalence at level ${\mathrm{n}}$ for all ${\mathrm{n}}>0$ and ${\mathcal{O}}\to\mathrm{Comm}_{\mathcal{C}}$ induces an equivalence at level 00. ∎
There is the following relationship between ${\mathcal{O}}$ and ${\mathcal{O}}^{\mathrm{nu}}$-algebras:
@@
    There is a canonical embedding $\mathrm{Alg}_{{\mathcal{O}}^{\mathrm{nu}}}({\mathcal{C}})\hookrightarrow\mathrm{Alg}_{{\mathcal{O}}}({\mathcal{C}})_{/{\mathbb{1}}}$ sending ${\mathrm{X}}$ to ${\mathrm{X}}\oplus{\mathbb{1}}.$
@@
    If ${\mathcal{C}}$ admits fibers, the embedding $\mathrm{Alg}_{{\mathcal{O}}^{\mathrm{nu}}}({\mathcal{C}})\hookrightarrow\mathrm{Alg}_{{\mathcal{O}}}({\mathcal{C}})_{/{\mathbb{1}}}$ admits a right adjoint that takes the fiber of the augmentation.
@@
    The embedding $\mathrm{Alg}_{{\mathcal{O}}^{\mathrm{nu}}}({\mathcal{C}})\hookrightarrow\mathrm{Alg}_{{\mathcal{O}}}({\mathcal{C}})_{/{\mathbb{1}}}$ is an equivalence if ${\mathcal{C}}$ is stable or idempotent complete and additive.
@@
The forgetful functor $\mathrm{Alg}_{{\mathcal{O}}}({\mathcal{C}})\to\mathrm{Alg}_{{\mathcal{O}}^{\mathrm{nu}}}({\mathcal{C}})$ admits a left adjoint ${\mathrm{F}}:={\mathcal{O}}\circ_{{\mathcal{O}}^{\mathrm{nu}}}(-).$ For every ${\mathcal{O}}^{\mathrm{nu}}$-algebra ${\mathrm{X}}$ in ${\mathcal{C}}$ the unit ${\mathrm{X}}\to{\mathrm{F}}({\mathrm{X}})$ of the adjunction and the unit ${\mathbb{1}}\to{\mathrm{F}}({\mathrm{X}})$ of the ${\mathcal{O}}$-algebra give rise to a morphism $\theta_{\mathrm{X}}:{\mathrm{X}}\oplus{\mathbb{1}}\to{\mathrm{F}}({\mathrm{X}})$ in ${\mathcal{C}},$ which determines a natural transformation $\theta:(-)\oplus{\mathbb{1}}\to{\mathrm{F}}$ of functors $\mathrm{Alg}_{{\mathcal{O}}^{\mathrm{nu}}}({\mathcal{C}})\to{\mathcal{C}}$. For every ${\mathrm{Y}}\in{\mathcal{C}}$ the morphism
$\theta_{{\mathcal{O}}^{\mathrm{nu}}\circ{\mathrm{Y}}}:({\mathcal{O}}^{\mathrm{nu}}\circ{\mathrm{Y}})\oplus{\mathbb{1}}\to{\mathrm{F}}({\mathcal{O}}^{\mathrm{nu}}\circ{\mathrm{Y}})\simeq{\mathcal{O}}\circ{\mathrm{Y}}$
is the canonical equivalence. Because source and target of $\theta$ preserve sifted colimits and every ${\mathcal{O}}^{\mathrm{nu}}$-algebra is a geometric realization of free ${\mathcal{O}}^{\mathrm{nu}}$-algebras, $\theta$ is an equivalence.
Since ${\mathcal{C}}$ is compatible with the initial object, ${\mathcal{O}}^{\mathrm{nu}}\circ(-):{\mathcal{C}}\to{\mathcal{C}}$ preserves the zero object and so $\mathrm{Alg}_{{\mathcal{O}}^{\mathrm{nu}}}({\mathcal{C}})$ admits a zero object that lies over the zero object of ${\mathcal{C}}.$ So ${\mathrm{F}}$ lifts to a functor $\bar{{\mathrm{F}}}:\mathrm{Alg}_{{\mathcal{O}}^{\mathrm{nu}}}({\mathcal{C}})\to\mathrm{Alg}_{{\mathcal{O}}}({\mathcal{C}})_{/{\mathbb{1}}}$ that factors through the full subcategory $\mathrm{Alg}_{{\mathcal{O}}}({\mathcal{C}})_{/{\mathbb{1}}}^{\prime}\subset\mathrm{Alg}_{{\mathcal{O}}}({\mathcal{C}})_{/{\mathbb{1}}}$ of augmented algebras whose augmentation admits a fiber in ${\mathcal{C}}.$

### v2

@@ 2.4. Truncating $\infty$-operads
Let $\mathcal{C}$} be a symmetric monoidal $\infty$-category compatible with the initial object that admits small colimits. The embedding $\kappa:\Sigma_{\geq{n}}\subset\Sigma_{\geq 1}$ induces a colocalization
$\kappa_{!}:{\mathrm{sSeq}}({\mathcal{C}})_{\geq{n}}\rightleftarrows{\mathrm{sSeq}}({\mathcal{C}})_{\geq 1}:\kappa^{\ast},$
where the right adjoint is restriction and the fully faithful left adjoint extends under degree n{n} by the initial object. By Lemma 2.43 the embedding $\kappa_{!}:{\mathrm{sSeq}}({\mathcal{C}})_{\geq{n}}\subset{\mathrm{sSeq}}({\mathcal{C}})_{\geq 1}$ is ${\mathrm{sSeq}}({\mathcal{C}})_{\geq 1}$-linear. Thus for every non-unital $\infty$-operad $\mathcal{O}$} in $\mathcal{C}$} we obtain a colocalization
${\mathrm{RMod}}_{{\mathcal{O}}}({\mathrm{sSeq}}({\mathcal{C}})_{\geq{n}})\rightleftarrows{\mathrm{RMod}}_{{\mathcal{O}}}({\mathrm{sSeq}}({\mathcal{C}})_{\geq 1}):\tau_{\geq{n}}.$
@@
Let $\mathcal{C}$} be a presentably symmetric monoidal $\infty$-category and $\mathcal{O}$} a non-unital $\infty$-operad in $\mathcal{C}$} such that ${\mathcal{O}}_{1}={\mathbb{1}}$. Let ${n}\geq 1$ and $\mathrm{X}$} a right $\mathcal{O}$}-module in $\mathcal{C}$} concentrated in degree n{n}. The space of right $\mathcal{O}$}-module structures on $\mathrm{X}$} is contractible.
@@
As ${\mathrm{sSeq}}({\mathcal{C}})_{\geq 1}$ is presentable, the small colimits preserving functor $(-)\circ{\mathrm{X}}:{\mathrm{sSeq}}({\mathcal{C}})_{\geq 1}\to{\mathrm{sSeq}}({\mathcal{C}})_{\geq 1}$ admits a right adjoint $\mathrm{X}$},-]. Thus the space of right $\mathcal{O}$}-module structures on $\mathrm{X}$} is equivalent to the space of maps of $\infty$-operads ${\mathcal{O}}\to[{\mathrm{X}},{\mathrm{X}}]$ in $\mathcal{C}$}.
For every ${k}\in\Sigma_{\geq 1}$ the functor ${\mathrm{sSeq}}({\mathcal{C}})_{\geq 1}\to{\mathcal{C}}$ evaluating at k{k} admits a left adjoint ${\mathrm{F}}_{k}$ sending $\mathrm{Z}$} to the symmetric sequence concentrated in degree k{k} with value $\Sigma_{k}\times{\mathrm{Z}}.$ Hence by adjointness we get the following: If k>1{k}>1, we find that
${\mathcal{C}}({\mathrm{Z}},[{\mathrm{X}},{\mathrm{X}}]_{k})\simeq{\mathrm{sSeq}}({\mathcal{C}})_{\geq 1}({\mathrm{F}}_{k}({\mathrm{Z}})\circ{\mathrm{X}},{\mathrm{X}})$
is contractible so that ${\mathrm{X}},{\mathrm{X}}]_{k}$ is the final object. Otherwise, we find that
${\mathcal{C}}({\mathrm{Z}},[{\mathrm{X}},{\mathrm{X}}]_{1})\simeq{\mathrm{sSeq}}({\mathcal{C}})_{\geq 1}({\mathrm{F}}_{1}({\mathrm{Z}})\circ{\mathrm{X}},{\mathrm{X}})\simeq{\mathcal{C}}^{\Sigma_{n}}({\mathrm{Z}}\otimes{\mathrm{X}}_{n},{\mathrm{X}}_{n})\simeq{\mathcal{C}}({\mathrm{Z}},[{\mathrm{X}}_{n},{\mathrm{X}}_{n}]^{\Sigma_{n}}),$
where ${\mathrm{X}}_{n},{\mathrm{X}}_{n}]$ is the internal hom in $\mathcal{C}$}, and so ${\mathrm{X}},{\mathrm{X}}]_{1}\simeq[{\mathrm{X}}_{n},{\mathrm{X}}_{n}]^{\Sigma_{n}}$. Using the localization
$(-)_{\leq 1}:{\mathrm{Op}}({\mathcal{C}})^{\mathrm{nun}}\to{\mathrm{Op}}({\mathcal{C}})^{\mathrm{nun}}_{\leq 1}=\mathrm{Alg}({\mathcal{C}})$
the space of maps of $\infty$-operads ${\mathcal{O}}\to[{\mathrm{X}},{\mathrm{X}}]$ in $\mathcal{C}$} identifies with the space of maps ${\mathcal{O}}_{1}\to[{\mathrm{X}}_{n},{\mathrm{X}}_{n}]^{\Sigma_{n}}$ of associative algebras in $\mathcal{C}$}. But by assumption ${\mathcal{O}}_{1}\simeq{\mathbb{1}}$ so that ${\mathcal{O}}_{1}$ is the initial associative algebra in $\mathcal{C}$}.
@@ 2.5. Hopf $\infty$-operads
Let $\mathcal{C}$} be a symmetric monoidal $\infty$-category. By [39, Proposition 3.2.4.3.] the $\infty$-category $\mathrm{Alg}_{{\mathbb{E}}_{\infty}}({\mathcal{C}})$ carries a canonical symmetric monoidal structure such that the forgetful functor $\mathrm{Alg}_{{\mathbb{E}}_{\infty}}({\mathcal{C}})\to{\mathcal{C}}$ is symmetric monoidal. So dually, the $\infty$-category
$\mathrm{co}\mathrm{Alg}_{{\mathbb{E}}_{\infty}}({\mathcal{C}}):=\mathrm{Alg}_{{\mathbb{E}}_{\infty}}({\mathcal{C}}^{\mathrm{op}})^{\mathrm{op}}$
carries a symmetric monoidal structure such that the forgetful functor $\mathrm{co}\mathrm{Alg}_{{\mathbb{E}}_{\infty}}({\mathcal{C}})\to{\mathcal{C}}$ is symmetric monoidal.
If $\mathcal{C}$} admits small colimits, $\mathrm{co}\mathrm{Alg}_{{\mathbb{E}}_{\infty}}({\mathcal{C}})$ admits small colimits and the forgetful functor $\mathrm{co}\mathrm{Alg}_{{\mathbb{E}}_{\infty}}({\mathcal{C}})\to{\mathcal{C}}$ preserves small colimits [39, Corollary 3.2.2.5.]. So we can make the following definition:
@@
A Hopf $\infty$-operad in $\mathcal{C}$} is an $\infty$-operad in $\mathrm{co}\mathrm{Alg}_{{\mathbb{E}}_{\infty}}({\mathcal{C}})$.
@@
${\mathrm{Op}}_{\mathrm{Hopf}}({\mathcal{C}}):={\mathrm{Op}}(\mathrm{co}\mathrm{Alg}_{{\mathbb{E}}_{\infty}}({\mathcal{C}})),\ {\mathrm{Op}}_{\mathrm{Hopf}}^{\mathrm{nun}}({\mathcal{C}}):={\mathrm{Op}}(\mathrm{co}\mathrm{Alg}_{{\mathbb{E}}_{\infty}}({\mathcal{C}}))^{\mathrm{nun}}.$
The symmetric monoidal forgetful functor $\mathrm{co}\mathrm{Alg}_{{\mathbb{E}}_{\infty}}({\mathcal{C}})\to{\mathcal{C}}$ yields forgetful functors
${\mathrm{Op}}_{\mathrm{Hopf}}({\mathcal{C}})\to{\mathrm{Op}}({\mathcal{C}}),\ {\mathrm{Op}}^{\mathrm{nun}}_{\mathrm{Hopf}}({\mathcal{C}})\to{\mathrm{Op}}({\mathcal{C}})^{\mathrm{nun}}.$
@@
Since $\mathrm{co}\mathrm{Alg}_{{\mathbb{E}}_{\infty}}({\mathcal{C}})$ admits a final object lying over the tensor unit of $\mathcal{C}$}, by Lemma 2.32 the $\infty$-categories ${\mathrm{Op}}_{\mathrm{Hopf}}({\mathcal{C}}),\ {\mathrm{Op}}^{\mathrm{nun}}_{\mathrm{Hopf}}({\mathcal{C}})$ admit a final object lying over the constant symmetric sequence (concentrated in positive degrees) with value the tensor unit.
@@
The non-unital commutative $\infty$-operad $\mathrm{Comm}^{\mathrm{nun}}_{\mathcal{C}}$ in $\mathcal{C}$} is the final object of ${\mathrm{Op}}^{{\mathrm{nun}}}_{\mathrm{Hopf}}({\mathcal{C}})$.
@@
The non-counital cocommutative $\infty$-cooperad ${\mathrm{coComm}}^{\mathrm{nun}}_{\mathcal{C}}$ in $\mathcal{C}$} is $\mathrm{Comm}^{\mathrm{nun}}_{{\mathcal{C}}^{\mathrm{op}}}$.
@@
${\mathcal{O}}^{\mathrm{nun}}:=\mathrm{Comm}_{\mathcal{C}}^{\mathrm{nun}}\times_{\mathrm{Comm}_{\mathcal{C}}}{\mathcal{O}}$
@@
$\mathrm{Comm}_{\mathcal{C}}^{\mathrm{nun}}\times_{\mathrm{Comm}_{\mathcal{C}}}{\mathcal{O}}$
in ${\mathrm{Op}}({\mathcal{C}})$ exists and is preserved by the forgetful functor ${\mathrm{Op}}({\mathcal{C}})\to{\mathrm{sSeq}}({\mathcal{C}})$ by Lemma 2.32. So it is enough to see that the fiber product $\mathrm{Comm}_{\mathcal{C}}^{\mathrm{nun}}\times_{\mathrm{Comm}_{\mathcal{C}}}{\mathcal{O}}$ in ${\mathrm{Hopf}}({\mathcal{C}})$ exists and is preserved by the forgetful functor ${\mathrm{Hopf}}({\mathcal{C}})\to{\mathrm{sSeq}}({\mathcal{C}})$. Since every fiber product in ${\mathrm{sSeq}}(\mathrm{co}\mathrm{Alg}_{{\mathbb{E}}_{\infty}}({\mathcal{C}}))$ of maps in ${\mathrm{Hopf}}({\mathcal{C}})$ lifts to a fiber product in ${\mathrm{Hopf}}({\mathcal{C}})$ by Lemma 2.32, it remains to prove that the fiber product $\mathrm{Comm}_{\mathcal{C}}^{\mathrm{nun}}\times_{\mathrm{Comm}_{\mathcal{C}}}{\mathcal{O}}$ in ${\mathrm{sSeq}}(\mathrm{co}\mathrm{Alg}_{{\mathbb{E}}_{\infty}}({\mathcal{C}}))$ exists and is preserved by the forgetful functor ${\mathrm{sSeq}}(\mathrm{co}\mathrm{Alg}_{{\mathbb{E}}_{\infty}}({\mathcal{C}}))\to{\mathrm{sSeq}}({\mathcal{C}})$. For that it is enough to show that for every ${n}\geq 0$ the fiber product $(\mathrm{Comm}_{\mathcal{C}}^{\mathrm{nun}})_{n}\times_{(\mathrm{Comm}_{\mathcal{C}})_{n}}{\mathcal{O}}_{n}$ in $\mathrm{co}\mathrm{Alg}_{{\mathbb{E}}_{\infty}}({\mathcal{C}})$ exists and is preserved by the forgetful functor $\mathrm{co}\mathrm{Alg}_{{\mathbb{E}}_{\infty}}({\mathcal{C}})\to{\mathcal{C}}$. This holds by [39, Proposition 3.1.3.3.] if and only if for every $\ell\geq 0$ the morphism
$((\mathrm{Comm}_{\mathcal{C}}^{\mathrm{nun}})_{n}\times_{(\mathrm{Comm}_{\mathcal{C}})_{n}}{\mathcal{O}}_{n})^{\otimes\ell}\to((\mathrm{Comm}_{\mathcal{C}}^{\mathrm{nun}})_{n})^{\otimes\ell}\times_{((\mathrm{Comm}_{\mathcal{C}})_{n})^{\otimes\ell}}({\mathcal{O}}_{n})^{\otimes\ell}$
is an equivalence. The latter map trivially is an equivalence since $\mathrm{Comm}_{\mathcal{C}}^{\mathrm{nun}}\to\mathrm{Comm}_{\mathcal{C}}$ induces an equivalence at level n{n} for all n>0{n}>0 and ${\mathcal{O}}\to\mathrm{Comm}_{\mathcal{C}}$ induces an equivalence at level 0. ∎
There is the following relationship between $\mathcal{O}$} and ${\mathcal{O}}^{\mathrm{nun}}$-algebras:
@@
    There is a canonical embedding $\mathrm{Alg}_{{\mathcal{O}}^{\mathrm{nun}}}({\mathcal{C}})\hookrightarrow\mathrm{Alg}_{{\mathcal{O}}}({\mathcal{C}})_{/{\mathbb{1}}}$ sending $\mathrm{X}$} to ${\mathrm{X}}\oplus{\mathbb{1}}.$
@@
    If $\mathcal{C}$} admits fibers, the embedding $\mathrm{Alg}_{{\mathcal{O}}^{\mathrm{nun}}}({\mathcal{C}})\hookrightarrow\mathrm{Alg}_{{\mathcal{O}}}({\mathcal{C}})_{/{\mathbb{1}}}$ admits a right adjoint that takes the fiber of the augmentation.
@@
    The embedding $\mathrm{Alg}_{{\mathcal{O}}^{\mathrm{nun}}}({\mathcal{C}})\hookrightarrow\mathrm{Alg}_{{\mathcal{O}}}({\mathcal{C}})_{/{\mathbb{1}}}$ is an equivalence if $\mathcal{C}$} is stable or idempotent complete and additive.
@@
The forgetful functor $\mathrm{Alg}_{{\mathcal{O}}}({\mathcal{C}})\to\mathrm{Alg}_{{\mathcal{O}}^{\mathrm{nun}}}({\mathcal{C}})$ admits a left adjoint ${\mathrm{F}}:={\mathcal{O}}\circ_{{\mathcal{O}}^{\mathrm{nu}}}(-).$ For every ${\mathcal{O}}^{\mathrm{nu}}$-algebra $\mathrm{X}$} in $\mathcal{C}$} the unit ${\mathrm{X}}\to{\mathrm{F}}({\mathrm{X}})$ of the adjunction and the unit ${\mathbb{1}}\to{\mathrm{F}}({\mathrm{X}})$ of the $\mathcal{O}$}-algebra give rise to a morphism $\theta_{\mathrm{X}}:{\mathrm{X}}\oplus{\mathbb{1}}\to{\mathrm{F}}({\mathrm{X}})$ in $\mathcal{C}$}, which determines a natural transformation $\theta:(-)\oplus{\mathbb{1}}\to{\mathrm{F}}$ of functors $\mathrm{Alg}_{{\mathcal{O}}^{\mathrm{nun}}}({\mathcal{C}})\to{\mathcal{C}}$. For every ${\mathrm{Y}}\in{\mathcal{C}}$ the morphism
$\theta_{{\mathcal{O}}^{\mathrm{nun}}\circ{\mathrm{Y}}}:({\mathcal{O}}^{\mathrm{nun}}\circ{\mathrm{Y}})\oplus{\mathbb{1}}\to{\mathrm{F}}({\mathcal{O}}^{\mathrm{nun}}\circ{\mathrm{Y}})\simeq{\mathcal{O}}\circ{\mathrm{Y}}$
is the canonical equivalence. Because source and target of $\theta$ preserve sifted colimits and every ${\mathcal{O}}^{\mathrm{nun}}$-algebra is a geometric realization of free ${\mathcal{O}}^{\mathrm{nun}}$-algebras, $\theta$ is an equivalence.
Since $\mathcal{C}$} is compatible with the initial object, ${\mathcal{O}}^{\mathrm{nun}}\circ(-):{\mathcal{C}}\to{\mathcal{C}}$ preserves the zero object and so $\mathrm{Alg}_{{\mathcal{O}}^{\mathrm{nun}}}({\mathcal{C}})$ admits a zero object that lies over the zero object of $\mathcal{C}$}. So $\mathrm{F}$} lifts to a functor $\bar{{\mathrm{F}}}:\mathrm{Alg}_{{\mathcal{O}}^{\mathrm{nun}}}({\mathcal{C}})\to\mathrm{Alg}_{{\mathcal{O}}}({\mathcal{C}})_{/{\mathbb{1}}}$ that factors through the full subcategory $\mathrm{Alg}_{{\mathcal{O}}}({\mathcal{C}})_{/{\mathbb{1}}}^{\prime}\subset\mathrm{Alg}_{{\mathcal{O}}}({\mathcal{C}})_{/{\mathbb{1}}}$ of augmented algebras whose augmentation admits a fiber in $\mathcal{C}$}.